OpenStudy (anonymous):

(MEDAL AND FAN PLUS A COOKIE GUYSSS :D♥♥) 7. -4 = g+1 (1 point) A. g= -4 B. g= 4 C. g= -5 D. g= -3

OpenStudy (anonymous):

use algebra. solve for g by subtracting -1 from both sides

OpenStudy (anonymous):

or subtracting 1 I should say

OpenStudy (anonymous):

o_O oki.

OpenStudy (anonymous):

\[g+1-1=-4-1\]

OpenStudy (yanasidlinskiy):

-4 = 1g + 1 -1 -1 -5 = 1g ------Divide by 1 -5 = g

OpenStudy (anonymous):

ohhhh okay.

OpenStudy (anonymous):

may I ask any more? ::P

OpenStudy (anonymous):

Division by 1 is unnecessary. But other than that, both @VeritasVosLiberabit and @YanaSidlinskiy are correct. Your answer is C :)

OpenStudy (anonymous):

ohh okay. thank you :) can I ask some more pretty please?

OpenStudy (yanasidlinskiy):

Yes. Go ahead:)

OpenStudy (anonymous):

yay :D thank you

OpenStudy (yanasidlinskiy):

I only got a few minutes. Hurry before I go...

OpenStudy (anonymous):

8. w- 2=-3 a. w= -3/2 b. w= -5 c. w= -1 dw= 6

OpenStudy (anonymous):

w - 2 = -3 w - 2 + 2 = -3 +2 w = -1 C :)

OpenStudy (yanasidlinskiy):

1w -2 = -3 +2 +2 1w = -1 Divide and you get -1...Sorry I'm kinda slow...

OpenStudy (anonymous):

it's okay. :) heehee. thank you guys so much! I really need the help cause I gotta hurry and get this work done before time runs out

OpenStudy (anonymous):

It's what we're here for :)

OpenStudy (yanasidlinskiy):

hahahahaha!!:) Ok...Gotta go now:) Glad I could help you:)

OpenStudy (anonymous):

oki thank you ^_^

OpenStudy (anonymous):

@girlnotonfire could you help me with some more ? :D

OpenStudy (yanasidlinskiy):

NO problem!!:)

OpenStudy (anonymous):

Sure! :)

OpenStudy (anonymous):

9. 6.4= 4x a. x= 2.4 b.x= 1.6 c. x= 4 d. x = 10.4

OpenStudy (anonymous):

6.4 = 4x 6.4/4 = 4x/4 <--- divide both sides by 4 1.6 = x B :)

OpenStudy (anonymous):

11. -7= (11/17)z a. z= -17/77 b. z= -77/17 c. z= -199/11 d. z= -11/119

OpenStudy (anonymous):

:) thank you again ^_^ you're the best

OpenStudy (anonymous):

Sorry, I wasn't getting notifications and my Internet is being all screwy >.< Anyway, -7 = (11/17)z -7/(11/17)=(11/17)z/(11/17) <--- divide both sides by 11/17 -119/11=z Hmmm is C supposed to say -119 instead of -199?